### v1

# The Spatial Complexity of Optical Computing and How to Reduce It

Abstract
Similar to algorithms, which consume time and memory to run, hardware requires resources to function. For devices processing physical waves, implementing operations needs sufficient “space,” as dictated by wave physics. How much space is needed to perform a certain function is a fundamental question in optics, with recent research addressing it for given mathematical operations, but not for more general computing tasks, e.g., classification. Inspired by computational complexity theory, we study the “spatial complexity” of optical computing systems in terms of scaling laws—specifically, how their physical dimensions must scale as the dimension of the mathematical operation increases—and propose a new paradigm for designing optical computing systems: space-efficient neuromorphic optics, based on structural sparsity constraints and neural pruning methods motivated by wave physics (notably, the concept of “overlapping nonlocality”). On two mainstream platforms, free-space optics and on-chip integrated photonics, our methods demonstrate substantial size reductions (to $1\%-10\%$ the size of conventional designs) with minimal compromise on performance. Our theoretical and computational results reveal a trend of diminishing returns on accuracy as structure dimensions increase, providing a new perspective for interpreting and approaching the ultimate limits of optical computing—a balanced trade-off between device size and accuracy.
The introduction of the Turing machine [1] marks a historic tipping point when humankind began to develop a deep mathematical understanding of computational machinery. Since then, when evaluating whether one problem is harder than another, we can rigorously examine the problems by analyzing their “computational complexity.” Understanding computational complexity has played an unrivaled role in modern computer science: The pursuit of solving (or approximating) as many problems as possible within polynomially constrained time and memory space not only promotes more efficient implementations but also deepens our understanding of the mathematical and physical mechanisms underlying the algorithms. In this Golden Age of AI, although computational resources have been increasing at an accelerating pace [2] and deep learning algorithms often prioritize performance over a transparently interpretable formalism [3, 4], studying computational complexity is still vital, as these investigations may reveal the scaling law(s) behind foundation models such as large language models (LLMs), thereby enabling more strategic resource allocation [5, 6] and shedding light on their remarkable performance [7, 8].
Over the past decades, the analysis of complexity has been extended to the studies of quantum systems [9, 10], nervous systems [11], and VLSI architecture [12], aiming to explore new science and optimize engineering solutions. On the other hand, sensing, imaging, and pre-processing hardware, which are often based on optical (or, more generally, wave-based) platforms, are usually overlooked in complexity analysis because they do not directly perform full-fledged computational tasks. Nevertheless, the increasing demand for low latency, high bandwidth/throughput, and energy-efficient data processing has promoted developments in optical pre-processing architectures for various applications, including autonomous driving (particularly with LiDARs [13]), augmented and virtual reality (AR/VR) [14], and high-speed biological microscopy [15]. In each application, light is engineered in a unique, objective-oriented way, and therefore, manipulating light to perform specific functionalities can indeed be considered a form of computation. Because one of the most critical limiting factors of hardware is its spatial dimensions [16], one could raise the following general question on the “computational complexity” of the hardware: Given a functionality to realize, how small can the optical hardware be? Or, in other words, what are the fundamental scaling laws that govern the size of a wave-based computing system as a function of the complexity of the task it is designed to perform?
Recently, D. A. B. Miller laid the foundation for these investigations by mapping the desired input-to-output function to a precisely defined concept of “overlap” between independent “communication channels” within the device, which in turn determines size requirements, such as the minimum possible thickness of imaging systems [17]. These fundamental limits, however, only strictly apply to well-defined, exact mathematical operations (for example, spatial differentiation, i.e., edge detection), whereas it is not clear how these size requirements would apply to more general, less exact computing functions, including inference tasks with tolerable levels of inaccuracy and performance degradation (e.g., classification). As an example, while it is clear how thin an optical system can be to perform a spatial derivative [17, 18], it is still unclear how thin it can be to perform image classification.
Figure 1: Reducing the spatial complexity of optical computing. Improving the scaling laws of free-space optics (a-c) and photonic chips (d-f) following the proposed space-efficient design paradigm. (a-c) Illustration of simplifying a generic free-space optical system. The thickness scaling law can be reduced to $\mathcal{O}(N^{1/2})$ when the $N\times N$ kernel matrix, representing the input-output function of the optical system, is designed to exhibit a “local sparse” form. Coloured lines represent the coupling coefficients between input and output ports (sampling points of the corresponding field profiles). Visually, a sparse matrix has significantly fewer couplings than a dense one, while in a local kernel matrix, all couplings are only slightly inclined from vertical, whereas nonlocal kernels lack this feature. (d-f) Illustration of simplifying a generic two-dimensional photonic chip, composed of a mesh of Mach–Zehnder interferometers (MZIs). The block-diagonalization, understood from a graph perspective, breaks the $N\times N$ kernel matrix to a linear number $N/N^{\prime}$ of small, decoupled complete bipartite graphs, $K_{N^{\prime},N^{\prime}}$, each of which requires $N^{\prime}(N^{\prime}-1)/2$ MZIs to realize. Therefore, when each block is sufficiently small, block-diagonalization reduces the total number of required MZIs from quadratic to quasi-linear.
To address these fundamental questions and advance one step closer toward the ultimate limits of wave-based computing, we propose and demonstrated the following two-pronged approach: 1. By analyzing the spatial scaling laws of optics, we discover what types of optics can scale optimally, namely, what physical properties and constraints make their spatial dimensions increase slowly as the dimension of the performed mathematical operation increases (Fig. 1). 2. Drawing inspiration from modern algorithm design, we propose a new paradigm for optical design, aimed at an optimal balance between performance and spatial resources, i.e., area or volume, through imposing specific constraints that improve the device kernel’s scaling. Two mainstream optical platforms, free-space and on-chip, are considered.
Specifically, we demonstrate that, to reduce the spatial complexity and improve the scaling, the device kernel should exhibit physics- and platform-specific structural sparsity: the “local sparse” form for three-dimensional free-space optics, which reduces the linear dependence of device thickness on the dimension of the mathematical operation to a square-root dependence, and the “block-diagonal” form for two-dimensional photonic chips, which reduces the number of required Mach-Zehnder interferometers from quadratic to quasi-linear.
These unique structural sparsity constraints, however, clearly limit the range of mathematical operations that can be realized, as these forms not only reduce the number of parameters but also constrain how the remaining parameters participate in the operation. Can the target task still be accomplished with high accuracy using only the required forms of operations? This question is beyond the scope of traditional optical design. Fortunately, artificial neural networks offer a promising solution, as they tend to exhibit considerable redundancy in their parameters [19], providing opportunities for pruning—reducing a large number of parameters in a network by selectively removing non-essential neurons and weights, with minimal impact on model accuracy. Pruning has proven effective across diverse neural network architectures, from the early-stage convolutional neural networks (CNNs) [20] and multilayer perceptrons [21] to modern, advanced CNNs [22, 23] and Transformers, which underpin LLMs [24, 25, 26].
By leveraging the neuromorphic architectures of optical neural networks (ONNs), we translate structural sparsity constraints, originating from wave physics, into optics-specific pruning methods and train local sparse ONNs and block-diagonal ONNs using this approach. We further explore the trade-off between the extent of pruning and the accuracy of these structurally sparse ONNs and identify space-efficient structures that significantly compactify the device (to $1\%-10\%$ the size of conventional designs), with negligible or only slight compromise on accuracy. Our design paradigm could potentially alleviate part of the computational cost by shifting some operations from electronic to optical processors without a significant increase in device size, thereby paving the way for the development of hybrid optical-electronic edge devices that optimally balance resource consumption and performance.
I Results
Figure 2: Standard optical nonlocality and overlapping nonlocality. (a) Schematic of a 1D free-space optical system. The communication cone of an output port, defined as the set of all couplings to it, is characterized by two parameters, the horizontal shift $d_{\text{shift}}$ and the spanning range $w_{\text{cone}}$. An ideally local optical device would have both $d_{\text{shift}}$ and $w_{\text{cone}}$ equal to zero. (b) The overlapping nonlocality (ONL) $C$ associated with a transverse aperture (or “cut”) is the number of communication cones intersecting it. For example, $C=5$ for the green cut and $C=6$ for the purple cut. (c) Demonstration of calculating $C$ for an arbitrary cut for a 2D free-space optical system. For a given cut, an output contributes a one to $C$ only if its communication cone intersects the cut. (d) Definition of the in-plane distance $d_{\parallel}$ traversed by an optical coupling. For a pair of coupled ports $(i,j)$, $d_{\parallel}$ is the distance between them, projected onto the output (or input) plane. This distance measures the nonlocality of each individual coupling.
Figure 3: Scaling laws of three types of optical device kernels. Scaling laws of the maximum ONL, $\max(C)$, with respect to the mathematical operation dimension, $N$, for the three considered types of device kernels: (a) trivial sparse, (b) row sparse, and (c) local sparse kernels. In (a), lines and shaded regions (inset) represent the average and one standard deviation of numerical simulation results. The standard deviation is negligibly small. In (b,c), darker lines represent theoretical values derived from Eq. 2, and lighter lines represent numerical simulation results (see Methods and Supplementary Note 1).
I.1 The Spatial Complexity of Free-Space Optics
To implement a nontrivial mathematical operation, an optical device has to be “nonlocal” — its output at any position should depend on the input across a range of positions, which ultimately implies the need for sufficient thickness to incorporate enough “channels” to communicate sideways within the device [17, 27]. The relation between the concept of nonlocality and thickness (and, more broadly, spatial complexity) is, however, a subtle one.
Consider the configuration shown in Fig. 2a: In a 1D optical structure, each output port (a sampling point for the output field) is coupled to a set of input ports (sampling points for the input field), spanning a range over the input plane. These couplings between pairs of points (input $j$ and output $i$) encode the function of the optical system, namely, they are the elements $D_{ij}$ of its kernel operator (in matrix form) $\mathbf{D}$. Two parameters can be used to describe how spatially extended, or “nonlocal”, the range of these couplings is: $d_{\text{shift}}$, the horizontal shift of the center of this range relative to the output, and $w_{\text{cone}}$, the width of the range. We then refer to the cone-shaped region (highlighted in orange in Fig. 2a) that encloses all the couplings for an arbitrary output $i$ as the communication cone of $i$. The parameters $d_{\text{shift}}$ and $w_{\text{cone}}$ represent different aspects of the conventional notion of nonlocality in optics and wave physics. Importantly, however, very large values for these parameters do not imply that the optical system require multiple “sideways channels” and, thus, a large thickness. As argued in [17], a single-mode optical fiber with multiple taps, for example, can have arbitrarily large nonlocality, as quantified here by $d_{\text{shift}}$ and $w_{\text{cone}}$, but would only require a finite (small) thickness to accommodate that single mode. It is a different aspect of “nonlocality” that is directly related to the size requirements of optics, the so-called overlapping nonlocality (ONL), which arises if the input position range for one output point overlaps with that for another output point, requiring multiple sideways channels to implement the desired function [17].
From a geometrical perspective (Fig. 2b), the ONL associated with a transverse aperture $S$ is the number $C$ of communication cones that must cross from one side of $S$ to the other to implement the desired kernel. We refer to the transverse aperture as a “cut,” as it divides the output (and input) plane into two parts. For a device operating at an effective wavelength of $\lambda_{0}/n$ (where $n$ is the maximum refractive index in the device), the cut must be large enough to support $C$ sideways channels to realize those crossing communication cones—one sideways channel for each cone. Using diffraction arguments (assuming the system only contains transparent, non-absorbing materials), one can then determine the lower bound on the device thickness $t$ from the maximum value of $C$ throughout the device [17]:
$t\geq\begin{cases}\max(C)\frac{\lambda_{0}}{2(1-\cos\theta)n}&\text{for 1D},\\ \max(C)\frac{1}{l_{\text{cut}}}\left[\frac{\lambda_{0}}{2(1-\cos\theta)n}\right]^{2}&\text{for 2D},\end{cases}$ (1)
where the parameter $\theta$ represents the maximum allowed ray angle inside the device. Eq. 1, also distinguishes between 1D and 2D cases, since in 1D, the size of the aperture is the thickness $t$, whereas in 2D, $t\equiv A/l_{\text{cut}}$, where $A$ is the area of the cut and $l_{\text{cut}}$ is the length of its projection to the output (or input) plane (Fig. 2c).
These theoretical results imply that the thickness required to perform an exact mathematical operation, described by a kernel $\mathbf{D}$, can only be reduced by decreasing the wavelength or increasing the refractive index, neither of which may be desirable or possible [28]. This is valid for any free-space optical (or wave-based) analog computing system, as the ONL $C$ is purely determined by the mathematical form of the desired functionality and not by the details of the optical implementation.
Here, instead, we approach the problem of minimizing the thickness, and therefore the size, of the system from a different perspective—by modifying the form of the mathematical operation so that the thickness scales slowly as the operation dimension increases, while maintaining high accuracy for a given task. To achieve this, we first analyze the scaling laws of the ONL, $C$, for operations under various constraints, and clarify the relationship between the ONL, sparsity, and standard optical locality.
Consider a device kernel $\mathbf{D}$ of dimension $N\times N$, with input and output ports arranged in grids on two parallel planes (see Supplementary Note 1). For the same $\mathbf{D}$, the number of required sideways channels, $C$, may vary across different choices of cuts. The most important quantity is therefore $\max(C)$, which is associated with the cut that is crossed by the largest number of sideways channels, and the entire device must be sufficiently thick to accommodate all the associated channels (see Supplementary Note 1 for the definitions of valid and invalid cuts). To analyze how the form of $\mathbf{D}$ affects $\max(C)$, we first study the worst-case scenario, where every output is coupled to all input points, resulting in a completely filled matrix $\mathbf{D}$. In this case, $C=N$, since the communication cone of every output intersects every cut.
Next, we examine whether increasing sparsity in $\mathbf{D}$ (by introducing zeros, i.e., removing couplings) reduces $\max(C)$. Our analysis and numerical experiments (see Supplementary Note 1 for the proof and simulation) show that trivial sparsity does not help, except when $\mathbf{D}$ becomes “extremely sparse,” as shown in Fig. 3a. However, an overly sparse kernel cannot encode enough free parameters, largely limiting its ability to realize useful optical functions. Therefore, trivially increasing sparsity is not a feasible way to reduce $\max(C)$.
One seemingly promising approach is to impose more drastic structural sparsity by removing a fraction of output ports, as this completely prevents those outputs from contributing to $C$. In matrix terms, removing outputs corresponds to setting rows of $\mathbf{D}$ to zero. With a fraction $\rho_{\text{row}}$ of rows remaining activated, $\max(C)$ scales as $\rho_{\text{row}}N$, still depending linearly on $N$ (Fig. 3b).
Aiming for a better scaling law, we then encourage locality when imposing sparsity by constructing matrices whose entries $D_{ij}$ are nonzero only when the in-plane distance $d_{\parallel}(i,j)$ (Fig. 2d) is below a predefined threshold $\max(d_{\parallel})$. This, in turn, constrains all outputs’ communication cones, $d_{\text{shift}}+w_{\text{cone}}/2\leq\max(d_{\parallel})$. Matrices constructed in this way display a distinct feature: When viewed in a graph layout, all couplings deviate only slightly from vertical (Fig. 1c). We refer to these matrices as local sparse matrices. While reducing the in-plane distance crossed by individual couplings does not directly reduce the ONL, if this distance is, on average, small for all couplings, then the number of independent communication cones crossing any transverse aperture should be reduced, hence reducing the ONL. Our further analysis and simulation indeed reveal that, with local sparse matrices, $\max(C)$ scales sublinearly as $\mathcal{O}(N^{1/2})$ (Fig. 3c). Eq. 2 summarizes the scaling laws for the expected values of the maximum ONL for the three considered types of matrices, corresponding to different forms of device kernel (see Supplementary Note 1 for details and proofs),
$\mathbb{E}[\max(C)]=\begin{cases}N&=\mathcal{O}\left(N\right)\text{ for trivial sparse matrices},\\ \rho_{\text{row}}N&=\mathcal{O}\left(N\right)\text{ for row sparse matrices},\\ 2\max(d_{\parallel})\cdot\left[\sqrt{2N}-\max(d_{\parallel})\right]&=\mathcal{O}\left(N^{1/2}\right)\text{ for local sparse matrices with small $\max(d_{\parallel})$}.\end{cases}$ (2)
Figure 4: Space-efficient computing with free-space optics. Schematic of (a) conventional, (b) row sparse, and (c) local sparse ONNs performing a classification task on the fashion-MNIST dataset. (d-f) Thicknesses of the three interlayer regions in conventional (top row), row sparse (middle row), and local sparse (bottom row) ONNs. To demonstrate the trade-off between model thickness and accuracy, each model type is pruned using five different pruning thresholds (see Methods). Error bars represent one standard deviation of thickness (along $y$-axis) and accuracy (along $x$-axis) across eight different random seeds. A local sparse ONN with a pruning threshold of $\tau=0.01$, compared to a conventional ONN with $\tau=0.05$, drastically reduces the thickness while only compromising accuracy by $3.6\%$ on the fashion-MNIST dataset (vertical orange arrows; see Supplementary Note 6).
Our findings imply that, while enforcing sparsity alone is not helpful, reducing the degree of standard optical nonlocality, quantified by $d_{\text{shift}}$ and $w_{\text{cone}}$ and bounded by $\max(d_{\parallel})$, also reduces the scaling of the maximum ONL, $\max(C)$, with the operation dimension. Therefore, reducing the spatial complexity of a free-space optical computing structure requires a pruning method motivated by wave physics and fundamentally rooted in the concept of optical nonlocality. These results promise compact, scalable optical realizations of large-scale computations, provided that the target task can still be accomplished with high accuracy, as discussed next.
I.2 Space-Efficient Computing with Free-Space Optics: Local Sparse ONNs
Unlike artificial neural networks stored and manipulated in computer memory, optical (or, any wave-based) neural networks exist in the physical world. Many properties of these networks, including their optical locality, stem from the Euclidean distance defined in the position space, thus demanding physics-motivated pruning techniques. To translate the structural sparsity constraints identified above into an optics-specific pruning method, we leverage the recently proposed brain inspired modular training (BIMT) [29], which offers a perfect routine for pruning physical neural networks targeting for local sparse patterns. BIMT focuses on spatial networks [30], where each neuron in the network is assigned a position, enabling the definition of physical distances. To enhance locality, BIMT penalizes nonlocal channels using an additional loss term, $\mathcal{L}_{\text{nonlocal}}$ (see Methods), and strategically swaps neurons and the associated weights across layers (except for the input). These features make BIMT an ideal method for designing local sparse optical neural networks (LSONNs), suitable for ultrathin free-space optical devices.
To quantitatively examine the trade-off between the optical device’s minimum thickness and the model’s inference accuracy, we train LSONNs and their conventional counterparts on three datasets, MNIST [31], fashion-MNIST [32], and Kuzushiji-MNIST [33]. Then, we apply weight pruning to the resulting networks by removing weights with absolute values below a threshold $\tau$. (Here, weight pruning is also applied to conventional networks because, otherwise, the network would be fully connected, corresponding to an extremely thick device.) Each network has $4$ bias-free layers with $\left[784,100,100,10\right]$ neurons. The neurons in the input and hidden layers are arranged in square grids in position space. The $10$ neurons in the output layer are arranged in a ring. Each layer, except the last one, is followed by a SiLU (Swish) activation function [34]. Once the network is trained, we calculate the minimum possible device thickness (i.e., the minimum distance between each two successive layers) for a free-space optical implementation, by calculating the maximum ONL and applying Eq. 1.
Fig. 4 summarizes the thickness-accuracy trade-offs of conventional, row sparse, and local sparse ONNs. Remarkably, across all three datasets, LSONNs achieve a thickness reduction of more than one order of magnitude in $t_{1,2}$ and reduce $t_{3}$ by approximately a factor of $3$ compared to their conventional and row sparse counterparts, with a slight or negligible drop in accuracy (e.g., only $3.6\%$ on the fashion-MNIST dataset). This thickness-accuracy trade-off can be interpreted as diminishing returns on accuracy: beyond a certain level of optical complexity and accuracy, additional spatial resources lead to only marginal improvements in accuracy. Additionally, we attribute this substantial reduction in thickness to a strategic allocation of the ONL across all transverse apertures: In LSONNs, large values of $C$ always pass through apertures with large in-plane length $l_{\text{cut}}$, while in conventional and row sparse ONNs, large values of $C$ often need to pass through narrow apertures (with small $l_{\text{cut}}$), causing “information bottlenecks” that require greater thickness to accommodate the necessary $C$ (see Supplementary Note 2).
In summary, we have demonstrated training LSONNs as an example of the proposed space-efficient paradigm for designing free-space optical computing structures that use an available volume optimally to perform an intended computation task: First, we have analyzed the scaling laws for the ONL, the key physical quantity associated with the spatial complexity of optical systems, and discovered that mathematical operations in the local sparse form can result in thin free-space optical structures. Then, by applying a physics-motivated pruning method that targets the required local sparse form, we have trained specialized ONNs that meet the desired criteria. This procedure is directly applicable to the designs of diffractive-surface- and metalens-based ONNs [35, 36], potentially enabling a new generation of scalable ultrathin free-space optical devices.
I.3 Space-Efficient Computing on Photonic Chips: Analysis and Design
Figure 5: Space-efficient computing on photonic chips. (a) Trade-off between model accuracy and the degree of block-diagonalization for different block-diagonal models. Shaded regions represent one standard deviation of accuracy across eight different random seeds. (b-e) Weights of block-diagonal and unpruned models. For each model, the number of free parameters, the involved unitary transformations (including both left and right unitary matrices, $U$ and $V^{T}$, from the singular value decomposition of all weight matrices), and the total number of required MZIs are listed on the right.
Optical networks on integrated photonic chips are typically based on arrays of photonic waveguides and modulators and meshes of Mach-Zehnder Interferometers (MZIs) performing arbitrary linear operations on the waveguide inputs [37, 38]. Optimizing spatial resources in this context, therefore, demands reducing the number of required MZIs, $n_{\text{MZI}}$. When realizing a device kernel matrix $\mathbf{D}$ of dimension $M\times N$, an MZI-based architecture requires $n_{\text{MZI}}=\left[M(M-1)+N(N-1)\right]/2$ components to implement two unitary transformations, $U(M)$ and $U(N)$ [37]. However, if the kernel matrix could be reduced to a “block-diagonal” form, the scaling of $n_{\text{MZI}}$ with both $M$ and $N$ would be reduced from quadratic to quasi-linear, as a block-diagonal matrix only comprises a linear number of small blocks with each block only requiring a few MZIs. This simplification of the photonic chip architecture becomes more apparent from a graph perspective: Block-diagonalization disassembles the huge bipartite connectivity graph into small, decoupled (disconnected) components, resulting in a significant loss of parameters and, more importantly, inevitably sacrificing inter-block connectivity (coupling), as shown in the connectivity graphs in Fig. 1d-f. Similar to free-space optics, the question is then whether a given operation can still be performed using a space-efficient architecture of this type, and what level of performance degradation can be expected.
While this block-diagonalization approach is not expected to be applicable to all computational tasks, we find that relevant tasks solvable by neural networks are amenable to block-diagonalization, leading to a substantial reduction in spatial complexity. To optimize the trade-off between $n_{\text{MZI}}$ and the ONNs’ performance, we propose a two-step pruning method. First, we include in the loss function an extra term, $\mathcal{L}_{\text{off-bloc-diag}}$, that penalizes all off-block-diagonal entries. No parameter is removed at this stage. Once the weights are trained, they are loaded into specially designed block-diagonal-structured (BDS) linear layers, where only entries in the diagonal blocks are registered as learnable parameters, while all other entries remain zero and are not updatable. The model with BDS linear layers is further trained on the same dataset to fully adapt to the block diagonal structure (see Methods).
To evaluate the trade-off between $n_{\text{MZI}}$ and the model’s inference accuracy, we train networks with several degrees of block-diagonalization on the three MNIST-like datasets used in the previous section [31, 32, 33]. Each network has $5$ bias-free layers with $\left[784,100,100,10,10\right]$ neurons. Each layer, except the last one, is followed by a SiLU activation function [34]. We consider six different block-diagonalized architectures, named from “b-d1” to “b-d6”, where $n_{\text{MZI}}$ varies from approximately $3,200$ to approximately $22,600$. Fig. 5 summarizes the trade-off between $n_{\text{MZI}}$ and accuracy. These six models share one important feature, the weight (size $10\times 10$) connecting the last two layers remains unpruned, which we find is crucial for maintaining satisfactory accuracy. (Results on models with a pruned last weight are included in Supplementary Note 6.) Remarkably, model “b-d4”, with its largest block being only $10\times 10$ in size, achieves accuracies of $93.24\pm 0.40\%$, $84.51\pm 0.16\%$, and $66.76\pm 0.25\%$ on the MNIST, fashion-MNIST, Kuzushiji-MNIST datasets, respectively. Implementing “b-d4” using the MZI-grid architecture already requires fewer than $4,200$ MZIs, compared to the unpruned implementation’s approximately $0.3$ million MZIs—a reduction of about $98.7\%$. We also expect that emergent photonic components, for example, inverse-designed, programmable multiport couplers [39, 40, 41], could enable an even more compact and flexibly tunable on-chip implementation of similar models based on small, decoupled blocks.
Finally, we extend this block-diagonalization approach to 1. develop block-circulant ONNs, whose weight matrices consist of rows of cyclically shifting blocks, leveraging light diffraction for space-efficient computing (e.g.,[42]); and 2. prune models optimized for edge AI applications, e.g., MobileNetV2 [43] for CIFAR-10 [44] classification. We block-diagonalize the fully connected (FC) layer of MobileNetV2 and, with a largest block size of $10\times 10$, reduce $n_{\text{MZI}}$ by $\sim 99\%$ (considering the photonic implementation of FC layers only), while the accuracy decreases only slightly from $92.97\pm 0.24\%$ to $92.41\pm 0.25\%$. (See Supplementary Note 4 for both models’ configurations and performance.) With these extensions, we envision that a hybrid architecture—employing digital implementations for convolution layers, nonlinear activations, shortcut connections, etc., and photonic analog components for the classifier based on block-diagonal (or block-circulant) FC layers—could facilitate FLOP-, memory-, and space-efficient computation acceleration in real-world edge applications.
II Discussion
For future optical systems to help addressing the growing challenge of large-scale, ultrafast, energy-efficient parallel data processing in a scalable way, we believe their spatial complexity should be carefully assessed to avoid trading off the benefits afforded by optics with impractical size requirements. Specifically, we have shown that, to create optical systems that scale optimally with the operation dimension, their kernel operators need to exhibit physics- and platform-specific structurally sparse patterns: the “local sparse” form for free-space optics and the “block-diagonal” form for photonic chips. Our space-efficient design paradigm offers engineering solutions aligned with significantly improved scaling laws, as demonstrated by examples of local sparse and block-diagonal ONNs. The space-accuracy trade-offs in these space-efficient ONNs reveal a trend of diminishing returns on accuracy: once the optics reaches a certain level of complexity and accuracy, further investments in spatial resources yield marginal accuracy increase. In fact, we have identified optimal ONNs that could achieve comparable accuracy to unpruned networks while consuming only $1\%-10\%$ of their spatial resources. In light of this, we interpret the ultimate limits of optical computing as a balance between device dimensions and performance, rather than merely high performance metrics. To achieve an optimal space-accuracy trade-off, adjusting the global connectivity topology of the optical system is as important as fine-tuning well-understood optical degrees of freedom, such as the refractive index contrast—if not more than.
Furthermore, space-efficient ONNs offer the first example of exploiting the full potential of neural pruning with optical hardware. Although the pruning-induced sparsity offers great flexibility and enables high compression rates for neural networks, leveraging this sparsity to achieve fast inference typically requires specialized software (e.g., cuSPARSE [45]) or hardware (e.g., EIE [46]) accelerators [47]. This is because the memory access patterns in sparse operations may not fully utilize GPUs’ parallel architecture. Consequently, while a sparse matrix has fewer effective operations, the overhead of managing sparsity and irregular memory access patterns can offset these benefits. This mismatch is an example of poor synergy between the model and the computational hardware. We believe that optical neuromorphic analog computing, with its high energy efficiency, high throughput, and low latency [48, 49], when fully optimized for space efficiency, could be a promising candidate for achieving an excellent software-hardware synergy and ultimately leading to superior computational acceleration.
We also note that the resources used in optical computing are not limited to spatial dimensions. Systems operating on synthetic dimensions, such as frequency [50, 51], angular momentum [51], and time [52], would follow different spatial scaling laws, but we expect that any reduction in spatial complexity would be accompanied by increased computational complexity in other dimensions (e.g., operation time scaling as $\mathcal{O}(N^{2})$) and other trade-offs (e.g., time-accuracy trade-offs). The study of multi-dimensional scaling laws for wave-based computing systems remains an interesting open area of research. At last, we note that the two mentioned optical/photonic platforms were neither originally intended for neuromorphic computing nor for emulating complex (brain-like) systems. As depicted in Fig. 1, the connectivity within these architectures is highly regular, resulting in a simple graph topology compared to complex networks, such as power grids, brain networks, or social networks [53, 54, 55]. These complex networks can exhibit various unique phenomena, such as the small-world effect [56], where each node can reach any other node through only a few intermediary “repeater” nodes, even if the network is extremely large, indicating that these networks are highly efficient in information exchange [57, 58]. Could more intricate connectivity with richer topology enable even more space-efficient optics and reduce the spatial complexity scaling law below polynomial growth? Highly scattering media [59] may be a candidate that could facilitate such complex connectivity in a wave-based system.
III Data and Code Availability
All datasets used in this research are publicly available. The code for generating and analyzing sparse, row sparse, and local sparse matrices is available at https://github.com/lyd5039/The-Spatial-Complexity-of-Optical-Computing/.
IV Competing Interests
Y. L. and F. M. are listed as inventors on a U.S. provisional patent application (Serial No. 63/718,499) on training and pruning methods for designing space-efficient free-space optics and photonic chips.